OpenStudy (anonymous):

please help guys! in such desperte need! Complete the proof with the statements and reasons below. Write the letter of the statement or reason in the box. Statments: 1) 2) 3) 4) 5) x=15 Reasons: 1) Given 2) 3) 4) 5) Please match these up guys, here are the things to chose from, these reasons, have to match up right with the statments.

OpenStudy (anonymous):

A. Reflexive Property B. Symmetric Property C. x=15 D. Transitive Property E. Addition Property of Equality F. Subtraction Property of Equality G. 3x-12=33 H. Multiplication Property of Equality I. Division Property of Equality J. Substitution Property K. Additive Identity L. 3x=45 M. Multiplicative Identity N. Commutative Property O. 3x+0=45 P. Associative Property Q. Distributive Property R. Given S.

OpenStudy (anonymous):

s. 3(x-4)=33

OpenStudy (anonymous):

can someone just please help me! I really need it:(

OpenStudy (anonymous):

S1) S S2) G S3) O S4) L S5) C those are the steps to isolate and find the value of x.

OpenStudy (anonymous):

so those are the stetments right?

OpenStudy (anonymous):

Yes As for the reasons 2 to 5, it's been too many years since I've had to justify those steps, so I don't remember the exact names of each property. They should be in your notes

OpenStudy (anonymous):

alright, could you give me the reasons?

OpenStudy (anonymous):

that match those statments u just gve me above?

OpenStudy (anonymous):

no, sorry, it's too far in my memory. You should check in your notes or manual, the definition of each property should be there and it'll help you associate the properties (reasons) with the statements

OpenStudy (anonymous):

alrighty thanks:)

OpenStudy (anonymous):

no problem!

jimthompson5910 (jim_thompson5910):

We have the following Statments: 1) 2) 3) 4) 5) x = 15 Reasons: 1) Given 2) 3) 4) 5) So what is given here?

jimthompson5910 (jim_thompson5910):

ie, what is the original problem?

OpenStudy (anonymous):

I think that what you gave me should be flipped around like x=15 should be number five because I think its a whole problem.

OpenStudy (anonymous):

The given is 3(x-4)=33 right?

jimthompson5910 (jim_thompson5910):

good so that goes in statement 1) like this Statments: 1) 3(x-4)=33 2) 3) 4) 5) x = 15 Reasons: 1) Given 2) 3) 4) 5)

jimthompson5910 (jim_thompson5910):

We have 2 options, we can either divide or distribute, but we're forced to distribute because there is no option that has x-4 = 11 (which would indicate division)

jimthompson5910 (jim_thompson5910):

so this means we then get Statments: 1) 3(x-4)=33 2) 3x - 12 = 33 3) 4) 5) x = 15 Reasons: 1) Given 2) Distributive Property 3) 4) 5)

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (anonymous):

fir statments wouldnt it be 3) 3x+0=45 4) 3x=45 right?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

Statements: 1) 3(x-4)=33 2) 3x - 12 = 33 3) 3x + 0 = 45 4) 3x = 45 5) x = 15 Reasons: 1) Given 2) Distributive Property 3) 4) 5)

jimthompson5910 (jim_thompson5910):

Now we just need 3 reasons for the last 3 statements

OpenStudy (anonymous):

and then for reasons it would be 3) Subtraction Property of Equality 4)Additive Identity 5)Division Property of Equality right?

jimthompson5910 (jim_thompson5910):

close, #3 is wrong though

jimthompson5910 (jim_thompson5910):

3x - 12 = 33 3x - 12+12 = 33+12 ... Add 12 to both sides using the additive property of equality 3x + 0 = 45

OpenStudy (anonymous):

I ment Addition Property of Equality

OpenStudy (anonymous):

sorry lol

jimthompson5910 (jim_thompson5910):

you got it Statements: 1) 3(x-4)=33 2) 3x - 12 = 33 3) 3x + 0 = 45 4) 3x = 45 5) x = 15 Reasons: 1) Given 2) Distributive Property 3) Addition Property of Equality 4) Additive Identity 5) Division Property of Equality

OpenStudy (anonymous):

yay!!!!! lol

jimthompson5910 (jim_thompson5910):

lol congrats

OpenStudy (anonymous):

okay i have something that you could correct, will you check? 1) 5(x + 3) + 9 = 59 Given 2) 5x + 15 + 9 = 59 Distributive property 3) 5x + 24 = 59 additions property of equality 4) 5x + 0 = 35 subtraction property of equality 5) 5x = 35 addative identity 6) x = 7 division property of equality r these right?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

3 is wrong

jimthompson5910 (jim_thompson5910):

everything else is correct though

OpenStudy (anonymous):

what would three be?

jimthompson5910 (jim_thompson5910):

Is there a choice that says "combine like terms"? That's the most accurate description of what is going on in #3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

here are the choices A. Reflexive Property B. Symmetric Property C. Transitive Property D. Addition Property of Equality E. Subtraction Property of Equality F. Multiplication Property of Equality G. Division Property of Equality H. Substitution Property I. Additive Identity J. Multiplicative Identity K. Commutative Property L. Associative Property M. Distributive Property N. Given

jimthompson5910 (jim_thompson5910):

well, I guess the substitution property fits because you replace 15+9 with 24 and you can legally do this because 15+9 = 24

jimthompson5910 (jim_thompson5910):

usually though, you'll just say "combine like terms" if you go from 15+9 to 24

OpenStudy (anonymous):

yea but wouldnt that be 4?

jimthompson5910 (jim_thompson5910):

4 is the subtraction (not substitution) property of equality

jimthompson5910 (jim_thompson5910):

so you have that right

OpenStudy (anonymous):

pooay so for three just put substitution in?

jimthompson5910 (jim_thompson5910):

yeah that's what I would do given these options

jimthompson5910 (jim_thompson5910):

personally, I'd say "combine like terms" since it fits better, but can't do that here

OpenStudy (anonymous):

okay thanks so much!

jimthompson5910 (jim_thompson5910):

yw